OpenStudy (anonymous):

Need help with finding slope! What is the slope of the line perpendicular to the equation 4x+3y=9

OpenStudy (anonymous):

3/4

OpenStudy (anonymous):

How did you get that

OpenStudy (anonymous):

@ramit.dour

OpenStudy (anonymous):

\[m _{1\times}m _{2}=-1\] when two slopes are perpendicular

OpenStudy (anonymous):

I was going to use y=mx+b ,but I dont know how to implement it

OpenStudy (anonymous):

in slope intercept equaation of line \[y =mx+c\]

OpenStudy (anonymous):

y=mx+c....?????

OpenStudy (anonymous):

m is slop of line c is y co ordinate of point of intersect of line with y axis

OpenStudy (anonymous):

I was given slope intercept formula

OpenStudy (anonymous):

b, c whatever u take , it doesn't matter . read definition

OpenStudy (anonymous):

but do you know how to use y=mx+b

OpenStudy (anonymous):



OpenStudy (anonymous):

slope of line is tan(x) where x is value of angle made by line from the positive x axis in anti clockwise direction.